OpenStudy (anonymous):

If the length of each side of a square is increased by 4 in., the area of the square becomes 9 times larger. Find the perimeter of the original square.

OpenStudy (anonymous):

From the story you need to find an equation, that's the first step.

OpenStudy (anonymous):

Yes, I agree, but how? I'm lost on how to set it up.

OpenStudy (anonymous):

The story says something about the area of the square when you increase the sides. Let's a call the lenght of a side x. Then that area is (x+4)^2, right?

OpenStudy (anonymous):

Yes, that's right. So do I then foil that? Would that be my first step?

OpenStudy (anonymous):

You're going to have to do that sooner or later, but first I would find the equation from the story, so (x+4)^2 = ?

OpenStudy (anonymous):

I don't know. It didn't give me what it equals. It only said the area of the square will then be 9 times larger.

OpenStudy (anonymous):

You can get the equation from that bit of information.

OpenStudy (anonymous):

Would it be 9x, then?

OpenStudy (anonymous):

Almost, we're talking about area...

OpenStudy (anonymous):

Remember x was the lenght of a side, so 9x would be 9 times the length of a side, but we want 9 times the area.

OpenStudy (anonymous):

I don't 9x squared is right, is it?

OpenStudy (anonymous):

Sorry... been working on algebra word problems for a few hours and am a little out of it, as far as thinking 100% goes.

OpenStudy (anonymous):

No problem, so 9 x^2 indeed So we get (x+4)^2=9x^2

OpenStudy (anonymous):

Then I will set it all equal to zero, right?

OpenStudy (anonymous):

Hold on, you mean (x+4)^2-9x^2=0 I suppose?

OpenStudy (anonymous):

Yes, sir, that's what I mean.

OpenStudy (anonymous):

Great, that's a good Idea.

OpenStudy (anonymous):

Can you solve for x?

OpenStudy (anonymous):

So far, I have -8(x-2)(x+1)=0.

OpenStudy (anonymous):

looks good.

OpenStudy (anonymous):

When setting it equal to 0, what happens to the -8? How does that affect the answer? Does that even make sense?

OpenStudy (anonymous):

I have x=2, -1, but I didn't use the -8 part.... so am I not done with the problem?

OpenStudy (anonymous):

It makes sense and it doesn't affect the answer: you can divide both sides by -8 to get (x-2)(x+1)=0

OpenStudy (anonymous):

So therefore, my answer is x=2 and x=-1?

OpenStudy (anonymous):

We're not done yet. The question was to compute the perimeter, not length of a side. Also are you sure both x=2 and x=-1 make sense? (You've solved the equation correctly)

OpenStudy (anonymous):

I won't use the x=-1. I'll cross it out, so to speak, and just work with the x=2, correct?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

So now I'll just plug my 2 in for any x in the original equation?

OpenStudy (anonymous):

You could do that, and see that it is in fact a solution to the equation, but I wouldn't do that here. You've found the length of one side, you want to find the perimeter of the square, so...

OpenStudy (anonymous):

The perimeter would be 8. Because 2x4=8, correct?

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

Thank you! :D

OpenStudy (anonymous):

Yay, you're welcome!

OpenStudy (anonymous):

Am I supposed to press "good answer" on the answers where you helped me? Does that give you more points or something? (I'm totally new to this, which I'm sure you could tell.)

OpenStudy (anonymous):

You could that, I'd get a medal. But I'm not exactly aching to get them :)

OpenStudy (anonymous):

After you get some medals you level up.

OpenStudy (anonymous):

:) Okay. Thank you again! I really appreciateed the help!